OpenStudy (anonymous):

What is negative 2 square root 2 squared?

OpenStudy (anonymous):

2 the negatives cancel and the root cancels

myininaya (myininaya):

\[-2\sqrt{(2)^2}=-2*2=-4\]

OpenStudy (anonymous):

I think its: (-2sqrt(2))^2=4(2)=8 Wording is bad.